Jennifer is 6 years older than her sister Janette. 4 years ago she was 4 times as old as her sister. how old is jennifer now?
Answer by macston(5194)

Jennifer is 6 years older than her sister Janette. 4 years ago she was 4 times as old as her sister. how old is jennifer now?
J=Jennifer's age=T+6; T=janette's age
J-4=4(T-4) Substitute for J
T+6-4=4(T-4)
T+2=4T-16 Add 16 to each side
T+18=4T Subtract T from each side.
18=3T Divide each side by 3
6=T ANSWER Janette is 6 years old.
T+6=6+6=12 ANSWER 2: Jennifer is 12 years old.
CHECK:
J-4 years=4(T-4 years)
12 years-4 years=4(6-4 years)
8 years=4(2 years)
8 years=8 years